# Explicit Performance Bound of Finite Blocklength Coded MIMO: Time-Domain versus Spatiotemporal Channel Coding

### v2

@@ III-B1 Achievability
$\log M_{\text{ST}}^{*}\left(n,\varepsilon\right)$ $\geq\log{\kappa_{\tau}^{n}\left(\mathcal{F}^{\text{ST}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)}-\log{\beta_{1-\varepsilon+\tau}^{n}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)}$
@@ III-B4 Maximal Achievable Rate
$\text{Var}\left[{{V}^{\text{ST}}}\left(\mathbb{H}\right)\right]\xrightarrow{{\rho}/{L}\;\gg 1}\xi\left[\frac{L\left(N-L\right)}{8N}+\frac{L\left(L-N\right)}{8N}\right]^{2}-\left[\frac{N\left(N-L\right)}{8L}+\frac{N\left(L-N\right)}{8L}\right]^{2}=0,$ (34)
@@ III-C Time-Domain Coding
In the case of time-domain coding, $L$ time-domain transmitted codewords ${{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}$ are utilized and their formations are mutually independent. The transmitted signal can be expressed as $\mathsf{X}=\left[{{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}\right]$. Given the requirement to apply the same codeword power constraints as in spatiotemporal coding, the codewords for time-domain coding need to be selected from the set ${{\mathcal{F}}^{\text{TD}}}=\left\{{{\mathbf{x}}_{i}}\in{{\mathbb{C}}^{n}}:{{\left\|{{\mathbf{x}}_{i}}\right\|}^{2}}=n\rho/L,i=1,\dots,L\right\}$. In this case, the information density of MIMO channel $\mathsf{H}$ can also be expressed as
$\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)$ $=\log\frac{d{{P}_{\mathbb{Y}|\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}}{d{{P}_{\mathbb{Y}|\mathbb{H}=\mathsf{H}}}}$
$=n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)$
$\quad+\text{tr}\left[\left({{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathsf{X}\mathsf{H}-\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}}\right)\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}\right]\log{e}$
@@
=$\overset{\left(b\right)}{\mathop{=}}\,n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)$
$\quad+\text{tr}\left[\left({{\mathsf{\Lambda}^{\text{H}}}{{{\tilde{\mathsf{X}}}}^{\text{H}}}\tilde{\mathsf{X}}\mathsf{\Lambda}+{{{\tilde{\mathbb{W}}}}^{\text{H}}}\tilde{\mathsf{X}}\mathsf{\Lambda}+{\mathsf{\Lambda}^{\text{H}}}{{{\tilde{\mathsf{X}}}}^{\text{H}}}\tilde{\mathbb{W}}-\frac{\rho}{L}\mathsf{\Lambda}^{2}{{{\tilde{\mathbb{W}}}}^{\text{H}}}\tilde{\mathbb{W}}}\right)\left({{{\mathsf{I}_{N}}+\frac{\rho}{L}\mathsf{\Lambda}^{2}}}\right)^{-1}\right]\log e$
@@
$=\sum\limits_{i=1}^{m}{\log\frac{d{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}}{d{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}}}$
$=\sum\limits_{i=1}^{m}{{{\iota}_{i}}\left({{{\mathbf{\tilde{x}}}}_{i}};{{{\mathbf{\tilde{Y}}}}_{i}}|{{\lambda}_{i}}\right)}.$ (38)
@@
$=n\log\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)+\text{tr}\left[\frac{{{\lambda}_{i}}\mathbf{\tilde{x}}_{i}^{\text{H}}{{{\mathbf{\tilde{x}}}}_{i}}+\sqrt{{{\lambda}_{i}}}{{{\mathbf{\tilde{W}}}}_{i}}^{\text{H}}{{{\mathbf{\tilde{x}}}}_{i}}+\sqrt{{{\lambda}_{i}}}\mathbf{\tilde{x}}_{i}^{\text{H}}{{{\mathbf{\tilde{W}}}}_{i}}-\frac{\rho}{L}{{\lambda}_{i}}{{{\mathbf{\tilde{W}}}}_{i}}^{\text{H}}{{{\mathbf{\tilde{W}}}}_{i}}}{1+\frac{\rho}{L}{{\lambda}_{i}}}\right]\log e.$ (39)
@@
${{\iota}_{i}}\left({{{\mathbf{\tilde{x}}}}_{i}};{{{\mathbf{\tilde{Y}}}}_{i}}|{{\lambda}_{i}}\right)=\sum\limits_{j=1}^{n}{\left[\log\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)+\left(1-\frac{{{\left|\sqrt{\frac{\rho}{L}{{\lambda}_{i}}}{{W}_{i,j}}-1\right|}^{2}}}{1+\frac{\rho}{L}{{\lambda}_{i}}}\right)\log e\right]}.$ (40)
In this case, according to Proposition 1, we can get that
$\beta_{1-\varepsilon,i}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)=-nC_{i}^{\text{TD}}\left(\mathsf{H}\right)+\sqrt{nV_{i}^{\text{TD}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)+\mathcal{O}\left(\log n\right).$ (41)
Therefore, in a channel use, the channel capacity and channel dispersion of link $i$ are respectively
@@ III-C1 Achievability
$\log M_{i}^{*}\left(n,\varepsilon\right)$ $\geq\log{\kappa_{\tau}^{n}\left(\mathcal{F}^{\text{TD}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}-\log{\beta_{1-\varepsilon+\tau}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}$
@@ III-C2 Converse
$\log M_{i}^{*}\left(n,\varepsilon\right)$ $\leq-\log{\beta_{1-\varepsilon}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}$
@@ V Simulation Results
Figure 2: Comparison of the achievability and converse bounds with the normal approximation in spatiotemporal coding with $L=N=4$, $\rho=0\text{ dB}$ and $\varepsilon=10^{-5}$.
Figure 3: Comparison of the achievability and converse bounds with the normal approximation in spatiotemporal coding with $L=N=4$, $\rho=10\text{ dB}$ and $\varepsilon=10^{-5}$.
Figure 4: Comparison of the achievability and converse bounds with the normal approximation in spatiotemporal coding with $L=N=8$, $\rho=10\text{ dB}$ and $\varepsilon=10^{-5}$.
We present Fig. 2, Fig. 3 and Fig. 4 in spatiotemporal coding as representatives for comparing the results of achievability (III-B1) and converse bounds (III-B2) with the normal approximation (31). The only difference between the parameter settings in Fig. 2 and Fig. 3 lies in the SNR, which is higher in Fig. 3. Moreover, the only difference in the parameter settings between Fig. 3 and Fig. 4 is in the spatial DoF, which is larger in Fig. 4. We employ the distance between the normal approximation and the converse bound to characterize the accuracy of the normal approximation. It is evidently discernible from the figures that the utilization of larger spatial DoF leads to a more accurate normal approximation, and the enhancement in accuracy is highly significant. The same holds true for the fact that an increment in SNR substantially improves the accuracy of the normal approximation. Consequently, it is rational to characterize the maximal achievable system performance via the normal approximation in our high SNR massive MIMO scenario.
@@ Appendix A Proof of Proposition 1
$\log{{\beta}^{n}_{\alpha}}\left({{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}},{{P}_{{\mathbf{Y}}}}\right)$ $\geq\frac{1}{\gamma}\left(\alpha-P\left[\frac{d{{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}}}{d{{P}_{{\mathbf{Y}}}}}\geq\gamma\right]\right)$
$=\frac{1}{\gamma}\left(\alpha-P\left[\log\frac{d{{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}}}{d{{P}_{{\mathbf{Y}}}}}\geq\log\gamma\right]\right)$
$=\frac{1}{\gamma}\left(\alpha-P\left[\iota\left({\mathbf{x}};{\mathbf{Y}}\right)\geq\log\gamma\right]\right).$ (58)
Let $T=\mathbb{E}\left[{{\left|\iota\left(x;Y\right)-C\right|}^{3}}\right]$, $B=6T/V^{\frac{3}{2}}$, $\Delta>0$ is a constant and $\log\gamma=nC+\sqrt{nV}\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)$, then we use the Berry-Esseen CLT and have
$P\left[\iota\left({\mathbf{x}};{\mathbf{Y}}\right)\geq\log\gamma\right]$ $=P\left[\frac{\iota\left({\mathbf{x}};{\mathbf{Y}}\right)-nC}{\sqrt{nV}}\geq\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)\right]$
$\leq\Phi\left(\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)\right)+\frac{B}{\sqrt{n}}$
$=\alpha-\frac{\Delta}{\sqrt{n}}.$ (59)
@@
$\log{{\beta}^{n}_{\alpha}}\left({{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}},{{P}_{{\mathbf{Y}}}}\right)$ $\geq-\log\gamma+\log\Delta-\frac{1}{2}\log n$
$=-nC-\sqrt{nV}\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)+\log\Delta-\frac{1}{2}\log n$
=$\overset{\left(a\right)}{\mathop{=}}-nC-\sqrt{nV}\Phi^{-1}\left(\alpha\right)+\mathcal{O}\left(\log n\right)$
=$\overset{\left(b\right)}{\mathop{=}}-nC+\sqrt{nV}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right),$ (60)
@@
Next, we let $\gamma_{0}=nC+\sqrt{nV}\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)$. Then, using the Berry-Esseen CLT in (11), we can have
$P\left[\iota\left({\mathbf{x}};{\mathbf{Y}}\right)\geq\log\gamma_{0}\right]$ $=P\left[\frac{\iota\left({\mathbf{x}};{\mathbf{Y}}\right)-nC}{\sqrt{nV}}\geq\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)\right]$
$\geq\Phi\left(\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)\right)-\frac{B}{\sqrt{n}}$
$=\alpha.$ (61)
@@
$\log{{\beta}_{\alpha}}\left({{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}},{{P}_{{\mathbf{Y}}}}\right)$ $\leq-\log\gamma_{0}$
$=-nC-\sqrt{nV}\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)$
=$\overset{\left(a\right)}{\mathop{=}}-nC-\sqrt{nV}\Phi^{-1}\left(\alpha\right)+\mathcal{O}\left(\log n\right)$
$=-nC+\sqrt{nV}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right),$ (62)

### v3

@@ III-B1 Achievability
$\log M_{\text{ST}}^{*}\left(n,\varepsilon\right)$
$\geq\log{\kappa_{\tau}^{n}\left(\mathcal{F}^{\text{ST}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)}-\log{\beta_{1-\varepsilon+\tau}^{n}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)}$
@@ III-B4 Maximal Achievable Rate
$\text{Var}\left[{{V}^{\text{ST}}}\left(\mathbb{H}\right)\right]\xrightarrow{{\rho}/{L}\;\gg 1}$ $\xi\left[\frac{L\left(N-L\right)}{8N}+\frac{L\left(L-N\right)}{8N}\right]^{2}$
$-\left[\frac{N\left(N-L\right)}{8L}+\frac{N\left(L-N\right)}{8L}\right]^{2}=0,$ (30)
@@ III-C Time-Domain Coding
In the case of time-domain coding, $L$ time-domain transmitted codewords ${{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}$ are utilized and their formations are mutually independent. The transmitted signal can be expressed as $\mathsf{X}=\left[{{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}\right]$. Given the requirement to apply the same codeword power constraints as in spatiotemporal coding, the codewords for time-domain coding need to be selected from the set ${{\mathcal{F}}^{\text{TD}}}=\left\{{{\mathbf{x}}_{i}}\in{{\mathbb{C}}^{n}}:{{\left\|{{\mathbf{x}}_{i}}\right\|}^{2}}=n\rho/L,i=1,\dots,L\right\}$. In this case, the information density of MIMO channel $\mathsf{H}$ can also be expressed as (III-C), which is shown in the botton of next page.
$\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)$ $=n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)+\text{tr}\left[\left({{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathsf{X}\mathsf{H}-\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}}\right)\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}\right]\log{e}$
@@
=$\overset{\left(b\right)}{\mathop{=}}\,n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)+\text{tr}\left[\left({{\mathsf{\Lambda}^{\text{H}}}{{{\tilde{\mathsf{X}}}}^{\text{H}}}\tilde{\mathsf{X}}\mathsf{\Lambda}+{{{\tilde{\mathbb{W}}}}^{\text{H}}}\tilde{\mathsf{X}}\mathsf{\Lambda}+{\mathsf{\Lambda}^{\text{H}}}{{{\tilde{\mathsf{X}}}}^{\text{H}}}\tilde{\mathbb{W}}-\frac{\rho}{L}\mathsf{\Lambda}^{2}{{{\tilde{\mathbb{W}}}}^{\text{H}}}\tilde{\mathbb{W}}}\right)\left({{{\mathsf{I}_{N}}+\frac{\rho}{L}\mathsf{\Lambda}^{2}}}\right)^{-1}\right]\log e$
@@
$=\sum\limits_{i=1}^{m}{\log\frac{d{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}}{d{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}}}=\sum\limits_{i=1}^{m}{{{\iota}_{i}}\left({{{\mathbf{\tilde{x}}}}_{i}};{{{\mathbf{\tilde{Y}}}}_{i}}|{{\lambda}_{i}}\right)}.$ (34)
@@
$=n\log\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)$
$\quad+\text{tr}\left[\frac{{{\lambda}_{i}}\mathbf{\tilde{x}}_{i}^{\text{H}}{{{\mathbf{\tilde{x}}}}_{i}}+\sqrt{{{\lambda}_{i}}}{{{\mathbf{\tilde{W}}}}_{i}}^{\text{H}}{{{\mathbf{\tilde{x}}}}_{i}}+\sqrt{{{\lambda}_{i}}}\mathbf{\tilde{x}}_{i}^{\text{H}}{{{\mathbf{\tilde{W}}}}_{i}}-\frac{\rho}{L}{{\lambda}_{i}}{{{\mathbf{\tilde{W}}}}_{i}}^{\text{H}}{{{\mathbf{\tilde{W}}}}_{i}}}{1+\frac{\rho}{L}{{\lambda}_{i}}}\right]\log e.$ (35)
@@
${{\iota}_{i}}\left({{{\mathbf{\tilde{x}}}}_{i}};{{{\mathbf{\tilde{Y}}}}_{i}}|{{\lambda}_{i}}\right)$
$\quad=\sum\limits_{j=1}^{n}{\left[\log\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)+\left(1-\frac{{{\left|\sqrt{\frac{\rho}{L}{{\lambda}_{i}}}{{W}_{i,j}}-1\right|}^{2}}}{1+\frac{\rho}{L}{{\lambda}_{i}}}\right)\log e\right]}.$ (36)
Proposition 2
Using the Berry-Esseen CLT, the Neyman-Pearson function for link $i$ in the MIMO channel under time-domain coding can be expressed as
$\beta_{1-\varepsilon,i}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)$
$\qquad=-nC_{i}^{\text{TD}}\left(\mathsf{H}\right)+\sqrt{nV_{i}^{\text{TD}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)+\mathcal{O}\left(\log n\right),$ (37)
where in a channel use, the channel capacity and channel dispersion of link $i$ are respectively
@@
Proof: The proof proceeds in the same way as Proposition 1, and can be considered as the case with one spatial DoF.
@@ III-C1 Achievability
$\log M_{i}^{*}\left(n,\varepsilon\right)$
$\geq\log{\kappa_{\tau}^{n}\left(\mathcal{F}^{\text{TD}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}$
$\quad-\log{\beta_{1-\varepsilon+\tau,i}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}$
@@ III-C2 Converse
$\log M_{i}^{*}\left(n,\varepsilon\right)$ $\leq-\log{\beta_{1-\varepsilon,i}^{n}\left({{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{{\mathbf{\tilde{X}}}}_{i}}={{{\mathbf{\tilde{x}}}}_{i}},{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}},{{P}_{{{{\mathbf{\tilde{Y}}}}_{i}}|{{\Lambda}_{i}}=\sqrt{{{\lambda}_{i}}}}}\right)}$
@@ V Simulation Results
(a) $L=N=4$, $\rho=0\text{ dB}$
(b) $L=N=4$, $\rho=10\text{ dB}$
(c) $L=N=16$, $\rho=10\text{ dB}$
Figure 2: Comparison of the achievability and converse bounds with the normal approximation in spatiotemporal coding with $\varepsilon=10^{-7}$.
@@
[…]
@@ Appendix A Proof of Proposition 1
$\log{{\beta}^{n}_{\alpha}}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)$
$\qquad\geq\frac{1}{\gamma}\left(\alpha-P\left[\frac{d{{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}}}{d{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}}\geq\gamma\right]\right)$
$\qquad=\frac{1}{\gamma}\left(\alpha-P\left[\log\frac{d{{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}}}{d{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}}\geq\log\gamma\right]\right)$
$\qquad=\frac{1}{\gamma}\left(\alpha-P\left[\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)\geq\log\gamma\right]\right).$ (54)
Since the capacity achieving input codewords are i.i.d, the input symbols at each channel use are independent of each other. Also, considering the assumption of a memoryless channel, we define $T=\mathbb{E}\left[{{\left|\frac{1}{n}\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)-{{C}^{\text{ST}}}\left(\mathsf{H}\right)\right|}^{3}}\right]$, $B=6T/\left({{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\right)^{\frac{3}{2}}$, $\Delta>0$ is a constant and $\log\gamma=n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)$, then we use the Berry-Esseen CLT and have
$P\left[\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)\geq\log\gamma\right]$
$\qquad=P\left[\frac{\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)}{\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}}\geq\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)\right]$
$\qquad\leq\Phi\left(\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)\right)+\frac{B}{\sqrt{n}}$
$\qquad=\alpha-\frac{\Delta}{\sqrt{n}}.$ (55)
@@
$\log{{\beta}^{n}_{\alpha}}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)$
$\qquad\geq-\log\gamma+\log\Delta-\frac{1}{2}\log n$
$\qquad=-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha-\frac{B+\Delta}{\sqrt{n}}\right)$
$\qquad\quad+\log\Delta-\frac{1}{2}\log n$
=$\qquad\overset{\left(a\right)}{\mathop{=}}-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha\right)+\mathcal{O}\left(\log n\right)$
=$\qquad\overset{\left(b\right)}{\mathop{=}}-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right),$ (56)
@@
Next, we let $\gamma_{0}=n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)$. Then, using the Berry-Esseen CLT in (10), we can have
$P\left[\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)\geq\log\gamma_{0}\right]$
$\qquad=P\left[\frac{\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)}{\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}}\geq\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)\right]$
$\qquad\geq\Phi\left(\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)\right)-\frac{B}{\sqrt{n}}$
$\qquad=\alpha.$ (57)
@@
$\log{{\beta}_{\alpha}}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)$
$\qquad\leq-\log\gamma_{0}$
$\qquad=-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha+\frac{B}{\sqrt{n}}\right)$
=$\qquad\overset{\left(a\right)}{\mathop{=}}-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\alpha\right)+\mathcal{O}\left(\log n\right)$
$\qquad=-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right),$ (58)